OpenStudy (anonymous):

With a tailwind, a plane flew 1650miles in 3 hrs. On the return flight against the same wind, the trip took 3 hours and 18 minutes. What is the speed of the plane? What is the speed of the wind?

OpenStudy (dumbcow):

d = rate*time Let s be speed of plane, w the speed of wind Note: 3 hr 18min = 3.3 hrs 1650 = 3(s+w) --> 550 = s+w 1650 = 3.3(s-w) --> 500 = s-w Add the equations 2s = 1050 s = 525 w = 550 - 525 = 25 Therefore, plane was going 525 mph, wind speed was 25mph

OpenStudy (anonymous):

Thank u! We just figured that out- but took a bit longer than u. :-)

OpenStudy (dumbcow):

:)